Suppose three coins are tossed. Let Y be the random variable representing the number of tails that occur. Find the values of the random variable Y

1
Expert's answer
2022-03-03T17:45:18-0500

Let's denote "H" - heads, "T" - tails.

There are "2^3=8" possible outcomes


"S=\{HHH, HHT, HTH, THH,"

"HTT, THT, TTH, TTT\}"

The possible values of the random variable "Y" are "0, 1, 2, 3."

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & Y \\ \hline
 HHH & 0\\
 \hdashline
 HHT & 1 \\
 \hdashline
 HTH & 1 \\
 \hdashline
 THH & 1 \\
 \hdashline
 HTT & 2 \\
 \hdashline
 THT & 2 \\
 \hdashline
 TTH & 2 \\
 \hdashline
 TTT & 3 \\
 \hdashline
 
\end{array}"


We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2"



Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 & 3 \\ \hline
 p(y) & 1/8 & 3/8 & 3/8 & 1/8 
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 & 3 \\ \hline
 p(y) & 0.125 & 0.375 & 0.375 & 0.125 
\end{array}"